Find the correlation coefficient and report it accurate to four decimal places.


A. Here is a bivariate data set.

xy
7347
8119
7611
7283
5586
6725
6476
6677
72-3
6255

Find the correlation coefficient and report it accurate to four decimal places.
r =

B.

The table below shows the number of state-registered automatic weapons and the murder rate for several Northwestern states.

xx11.587.13.72.52.72.30.5
yy13.71110.17.266.56.24.4

xx = thousands of automatic weapons
yy = murders per 100,000 residents

This data can be modeled by the equation y=0.85x+4.09.y=0.85x+4.09. Use this equation to answer the following;
Special Note: I suggest you verify this equation by performing linear regression on your calculator.

a) How many murders per 100,000 residents can be expected in a state with 4.1 thousand automatic weapons?

Answer =  Round to 3 decimal places.

b) How many murders per 100,000 residents can be expected in a state with 1.7 thousand automatic weapons?

Answer =  Round to 3 decimal places.

C.

A regression analysis was performed to determine if there is a relationship between hours of TV watched per day (xx) and number of sit ups a person can do (yy).

The results of the regression were:

y=ax+b a=-1.314 b=34.695 r2=0.603729 r=-0.777 Use this to predict the number of sit ups a person who watches 1.5 hours of TV can do, and please round your answer to a whole number.
Answer #1
X Y (x-x̅)² (y-ȳ)² (x-x̅)(y-ȳ)
73 47 17.64 0.360 -2.52
81 19 148.84 817.960 -348.92
76 11 51.84 1339.560 -263.52
72 83 10.24 1253.160 113.28
55 86 190.44 1474.560 -529.92
67 25 3.24 510.760 40.68
64 76 23.04 806.560 -136.32
66 77 7.8400 864.3600 -82.3200
72 -3 10.240 2560.360 -161.920
62 55 46.240 54.760 -50.320
ΣX ΣY Σ(x-x̅)² Σ(y-ȳ)² Σ(x-x̅)(y-ȳ)
total sum 688 476 509.600 9682.40 -1421.8
mean 68.80 47.60 SSxx SSyy SSxy

correlation coefficient ,    r = Sxy/√(Sx.Sy) = -1421.8/√(509.6*9682.4) = -0.6401

--------------

B)

a)  
y=0.85x+4.09  
x=4.1  
y=0.85*4.1+4.09=   7.575
  
b)  
y=0.85x+4.09  
x=1.7  
y=0.85*1.7+4.09=   5.535
  
  
C  
y=-1.314*x + 34.695  
x=1.5  
y=-1.314*1.5 + 34.695=   33